# Taming the Tail in Class-Conditional GANs: Knowledge Sharing via Unconditional Training at Lower Resolutions

### v1

@@ 4.1 Setup
Datasets. We use 6 different long-tailed datasets in our experiments: CIFAR10 and CIFAR100[19], LSUN [49], Flowers [31], iNaturalist2019 [11], and AnimalFaces [43]. This selection is intended to encompass a wide spectrum of image domains, dataset sizes, resolutions, and imbalance ratios (denoted as $\rho$). $\rho$ represents the ratio of the number of training samples between the classes with the most and the least number of examples. To ensure the tail classes stay few-shot, we keep the number of images in the smallest tail classes under 50 [47]. A detailed description of these datasets is provided in Supp. A.
@@
For more comprehensive details on the baselines, hyperparameter selections, and further implementation details, see Supp. F.
@@ 4.2 Evaluation Metrics For Long-Tail Datasets
First, when a larger and (more) balanced dataset following the same distribution as the long-tail training data is available, we use it to calculate metrics, e.g., full CIFAR10 before artificial imbalance. For naturally imbalanced datasets such as iNaturalist2019, we use the unmodified training set to calculate metrics. Note that when generating images for evaluation, we sample from the same distribution as the available real dataset, which may be imbalanced. While FID and KID indicate how well the real data distribution matches the generated images across all classes, we suggest additional metrics to mitigate the disproportionate influence of head classes in data statistic calculations.
In response to the varied number of training instances across classes in a long-tail setup, we propose evaluating metrics specifically on few-shot (FS) categories  [47]. We calculate the FID and KID for the few-shot subset and refer to them as “FID-FS” and “KID-FS”, respectively. This is tailored to evaluate the quality of generated samples in the tail classes. Contrasting with the standard FID/KID, we maintain an equal number of real images across all classes during our FID-FS/KID-FS calculation. This emphasizes learning quality on tail classes, irrespective of any imbalances that might be present in the few-shot subset. Reporting both standard and few-shot metrics provides a more comprehensive evaluation of long-tail learning, considering the performance of both head and tail classes.
@@ 4.3 Results
Below, we report the results obtained across different benchmarks and cGAN architectures. We pick the best model for reporting results as the one with the lowest FID-FS from two independent runs. It is important to note that our algorithm exhibits significantly less reliance on early stopping compared to baseline models, as illustrated in Fig. 2.
@@
Ablation on the Choice of Low-resolution for Unconditional Training ($\bm{res}_{\text{uc}}$). In our proposed method, one of the hyperparameter choices is to select an intermediate low-resolution $res_{\text{uc}}$ for unconditional training. All layers with equal or lower resolution than $res_{\text{uc}}$ do not receive class-conditional information. An unconditional GAN objective is added over the images and/or features at $res_{\text{uc}}$. To study the impact of $res_{\text{uc}}$, we conducted an ablation study on the AnimalFaces-LT dataset, containing images at $64\times 64$. Table. 2 presents the results for selecting $res_{\text{uc}}$ from resolutions lower than $64\times 64$, namely $8\times 8$, $16\times 16$, and $32\times 32$. Studying the obtained results, we observe that resolutions of $8\times 8$ and $16$ achieve relatively close performance. Conversely, the performance diminishes when layers up to $32\times 32$ are trained unconditionally. This is anticipated as the AnimalFaces-LT is at $64\times 64$, leaving only one up-sampling layer to learn the class-conditional information which is shown to be insufficient. In all our experiments, we use $res_{\text{uc}}$ of $8\times 8$ unless stated otherwise. Additional visual analysis on the role of $res_{\text{uc}}$ as well as an ablation on the choice of unconditional training objective weight $\lambda$ (see Eq. 2) are presented in the Supp. C.
@@
Can any low-res be the starting point for high-res generations? In general, our observations indicate that high-res images are guided, yet not entirely restricted by the low-res images, as shown in Fig. 4. Each row in this figure presents a set of different high-res images that are generated from the same unconditional low-res image. This illustration reveals that high-res images can substantially differ from low-resolution ones, both in terms of color and texture. This suggests that subsequent conditional blocks are capable of significant modification in the background or foreground.
Importance of FS metrics: comparison against unconditional training We also include results from training an unconditional StyleGAN2-ADA model [16] in Table 1, referred to as StyleGAN2-ADA UnCond. The unconditional model generates samples that follow the training distribution, which is mainly dominated by head classes. This bias favors unconditional learning in terms of the FID/KID metrics, which do not consider the skewness in the data distribution. Nevertheless, it can be observed the UnCond. baselines significantly perform worse in terms of few-shot metrics FID-FS/KID-FS. This suggests solely relying on FID/KID can be misleading. We strongly recommend including few-shot metrics, i.e., FID-FS and KID-FS, when evaluating cGANs on long-tailed datasets.
CIFAR10-LT In this benchmark, our method is compared with various baselines with imbalanced ratios $\rho$ of $50$ and $100$. This indicates the least frequent tail class includes only $100$ and $50$ training instances, respectively. Table. 3 quantitatively assesses the quality of the samples generated by our method and the baselines. All the baselines use the StyleGAN2-ADA [16] architecture. Our proposed method consistently outperforms the baselines across all metrics. This demonstrates that beyond improving generative learning across all classes (FID & KID), UTLO can also notably improve the learning from tail classes (FID-FS & KID-FS).
@@
Knowledge-sharing at Low Resolutions. As explained in Sec. 3, our proposed generator builds on top of low-resolution (e.g., $8\times 8$) unconditionally trained images which are subsequently used to generate conditional images at higher resolutions (e.g., $32\times 32$). Several low-resolution images are shown in Fig. 6 along with their high-res conditional images generated for the head and tail classes of the dataset. It can be seen that conditional images generated from the tail classes evidently share certain features from the head classes. This demonstrates that knowledge sharing at low resolutions is an effective approach for infusing information from head classes to tail ones. A quantitative analysis of knowledge-sharing is presented in Supp. D. For additional qualitative examples see Supp. H.
Figure 6: Knowledge sharing from head to tail classes in CIFAR10-LT dataset ($\rho=100$) using our proposed UTLO framework. The conditional images generated from the head (middle columns) and tail (right columns) classes share and are built on top of the same low-resolution (unconditional) images (left columns). Low-resolution images ($8\times 8$) are upsampled to that of CIFAR10-LT ($32\times 32$) for improved visualization.
CIFAR100-LT In addition to CIFAR10-LT, we evaluate our method and various baselines on the more challenging CIFAR100-LT dataset, wherein the tail classes include as few as five training instances. The quantitative results shown in Table. 4 reveal that our method outperforms the baselines in a benchmark where a large number of classes with high diversity are present. In Fig. 2, we demonstrate the FID-FS curve during the course of the training along with the generated samples from the tail classes of the CIFAR100-LT dataset. This shows the effectiveness of our proposed method in addressing mode collapse while regularization methods diverge and require early stopping. Further visual comparisons of our proposed method against the baselines for CIFAR100-LT can be found in Supp. H.
@@
LSUN5-LT We include LSUN5-LT as a higher resolution ($128\times 128$) and highly-imbalanced ($\rho=1000$) benchmark. In addition, we use Projected GAN with FastGAN[25] generator and DA [39]) as the baseline model. This is a different generator and discriminator design compared to StyleGAN2 along with a different augmentation method. In conditional FastGAN implementation, the class-specific information is injected using class-conditional batch normalization, differing from the style mapping network utilized in StyleGAN2. UTLO can be readily adapted to this design by using standard batch normalization layers at low resolutions and utilizing class-conditional batch normalization only at higher resolutions. However, it is noteworthy that methods such as [36] are restricted to StyleGANs and cannot be applied in this context.
Table. 5 compares the results obtained from our method against the baseline and with the addition of GSR [35]. Our method consistently surpasses the baselines by around $30\%$ across all metrics. As previously shown for the CIFAR10-LT (Table 3), UTLO not only improved on the tail classes but also significantly improved the dataset-wide FID/KID due to the high imbalance ratio ($\rho=1000$).
@@ 5 Conclusion
In this paper, we proposed UTLO, a novel framework designed to improve the training of cGANs on long-tailed data. Inspired by the observation that head and tail classes often have more similarities at lower resolutions, our method facilitates knowledge sharing from head to tail classes using unconditional training at lower resolutions. The proposed method enriches the limited training distribution of the tail classes and effectively addresses mode collapse, leading to significant improvement in image generation for tail classes. We have also introduced the FID-FS/KID-FS metric, an adaptation of widely-used GAN metrics, specifically tailored for tail classes. We hope our findings are useful for future work in long-tail learning.
@@ Appendix A Datasets
  • CIFAR10/100: The original CIFAR10 and CIFAR100 datasets consist of 50,000 training images at $32\times 32$ resolution with an equal number of images present in 10 and 100 classes, respectively. An imbalanced version of these datasets referred to as CIFAR10/100-LT, is widely used in the long-tail recognition literature [47]. We follow [5] in order to generate an exponentially long-tailed version with an imbalance ratio $\rho=\{50,100\}$ for CIFAR10-LT and $\rho=100$ for CIFAR100-LT. When calculating few-shot metrics, we follow [4] and take classes 6-9 in CIFAR10-LT and 70-99 in CIFAR100-LT as the few-shot subset in our evaluation.
@@
Across all datasets, we sort image categories in the decreasing order of their size, with class 0 containing the most training images and so on.
@@ Appendix B Effects of Weighted Sampling
where $n_{c}$ is the number of samples in class $c$, $\rho=\max_{c}\{n_{c}\}/\min_{c}\{n_{c}\}$ is the imbalance ratio, and $\beta\in[0,1]$ is a hyperparameter that damps the sampling imbalance. We illustrate this in Fig. 7 for CIFAR100-LT with $\rho=100$. When $\beta=0$, the sampling follows the original data distribution. As $\beta$ increases, the effective imbalance ratio $\varrho$ between the classes decreases. At $\beta=1$, head and tail classes have the same probability of being drawn during training, i.e. $\varrho=1$. While WS might help with more calibrated training, it does not promote knowledge sharing among classes.
Figure 7: WS CIFAR100-LT $(\rho=100)$. When $\beta=0$, the sampling follows the original data distribution. As $\beta$ increases, the effective imbalance ratio $\varrho$ between the classes decreases. At $\beta=1$, head and tail classes have the same probability of being drawn during training, i.e. $\varrho=1$.
Table. 6 presents a comparison of results obtained from the StyleGAN2-ADA baseline and the use of weighted sampling (WS). Our experiments indicate that WS improves the FID and KID metrics compared to the baseline. However, it does not enhance tail performance in terms of FID-FS and KID-FS, indicating that balancing alone is insufficient when there are very limited training examples in the tail. We also noticed that adding weighted sampling will boost overfitting. Fig. 8 illustrates the training FID-FS curves for the WS methods with different $\beta$ values, compared against our method and the baseline. For both CIFAR10-LT ($\rho=100$) and CIFAR100-LT ($\rho=100$) datasets, WS methods exhibit overfitting behavior. As $\beta$ increases, this becomes more pronounced, and training becomes unstable when $\beta=1$ across both datasets. While WS did not show any improvements for CIFAR100-LT, we found it to demonstrate relatively better performance than the baseline at the early stages of training before overfitting sets in. While our experiments on the role of WS in training cGANs in the long-tail setup are informative, we believe that reaching a comprehensive conclusion requires further analysis.
@@ Appendix C Ablation Study
Studying the results obtained in Table. 7, we oberve that resolutions of $8\times 8$ and $16$ achieve relatively close performance. On the other hand, the performance degrades when layers up to $32\times 32$ are trained unconditionally. This is anticipated as the AnimalFaces-LT is at $64\times 64$, leaving only one layer to learn the class-conditional information which is shown to be insufficient.
To better understand the role of intermediate low resolution selected during training, we show the unconditional low-res images $\hat{{\bm{x}}}_{l}$ from the intermediate layers (see Fig.3 in main paper) for different ablated resolutions in Fig. 9. The low-res images are upsampled to the same size for better visual comparison. As $res{\text{uc}}$ increases, more details are introduced to the unconditionally trained intermediate images. For the AnimalFaces-LT dataset, resolution $32\times 32$ already includes fine and definite details, making it challenging to generate (tail) class-specific changes in a single layer before reaching the output resolution of $64\times 64$. We demonstrate this in the bottom of Fig. 9 the final class-conditional images at $64\times 64$ resolution from the tail class bear are shown. When $res_{\text{uc}}$ is higher, the final output changes minimally from the lower-resolution images $\hat{{\bm{x}}}l$. On the other hand, at lower $res{\text{uc}}$, the level of change is higher.

### v2

@@ 4.1 Setup
Datasets. We use 6 different long-tailed datasets in our experiments: CIFAR10 and CIFAR100[20], LSUN [51], Flowers [32], iNaturalist2019 [11], and AnimalFaces [44]. This selection is intended to cover a wide spectrum of image domains, dataset sizes, resolutions, and imbalance ratios (denoted as $\rho$). $\rho$ represents the ratio of the number of training samples between the classes with the most and the least number of examples. To ensure the tail classes stay few-shot, we keep the number of images in the smallest tail classes under 50 [49]. A detailed description of these datasets is provided in Supp. A.
@@
For comprehensive details on the baselines, hyperparameters, and implementation, please refer to Supp. F.
@@ 4.2 Evaluation Metrics For Long-Tail Datasets
First, when a larger and (more) balanced dataset following the long-tail training data distribution is available, it is used for metric calculation, e.g., full CIFAR10 before artificial imbalance. For naturally imbalanced datasets such as iNaturalist2019, we use the unmodified training set to calculate metrics. Note that when generating images for evaluation, we sample from the same distribution as the available real dataset, which may be imbalanced. While FID and KID indicate how well the real data distribution matches the generated images across all classes, we suggest additional metrics to mitigate the disproportional influence of head classes in data statistic calculations.
In response to the varied number of training instances across classes in a long-tail setup, we propose evaluating metrics specifically on few-shot (FS) categories  [49]. We calculate the FID and KID for the few-shot subset and refer to them as “FID-FS” and “KID-FS”, respectively. This is tailored to evaluate the quality of the generated samples in the tail classes. Unlike the standard FID/KID, we maintain an equal number of real images across all classes during our FID-FS/KID-FS calculation. This emphasizes learning quality on tail classes, irrespective of any imbalances that might be present in the few-shot subset. Reporting both standard and few-shot metrics provides a more comprehensive evaluation of long-tail learning, considering the performance of both head and tail classes.
@@ 4.3 Results
In the following, we report the results obtained across different benchmarks and cGAN architectures. We pick the best model for reporting results as the one with the lowest FID-FS from two independent runs. Note that our method exhibits significantly less reliance on early stopping compared to the baselines, as illustrated in Fig. 2.
@@
Ablation on the Choice of Low-resolution for Unconditional Training ($\bm{res}_{\text{uc}}$). In our proposed method, one of the hyperparameter choices is to select an intermediate low-resolution $res_{\text{uc}}$ for unconditional training. All layers with equal or lower resolution than $res_{\text{uc}}$ do not receive class-conditional information. An unconditional GAN objective is added over the images and/or features at $res_{\text{uc}}$. To study the impact of $res_{\text{uc}}$, we conducted an ablation study on the AnimalFaces-LT dataset, containing images at $64\times 64$. Table. 2 presents the results for selecting $res_{\text{uc}}$ from resolutions lower than $64\times 64$, namely $8\times 8$, $16\times 16$, and $32\times 32$. When studying the obtained results, we observe that resolutions of $8\times 8$ and $16\times 16$ achieve relatively close performance. Conversely, the performance diminishes when layers up to $32\times 32$ are trained unconditionally. This is anticipated as AnimalFaces-LT is at $64\times 64$, leaving only one up-sampling layer to learn the class-conditional information which is shown to be insufficient. In all of our experiments, we use $res_{\text{uc}}$ of $8\times 8$ unless otherwise stated. Additional visual analysis on the role of $res_{\text{uc}}$ as well as an ablation on the choice of unconditional training objective weight $\lambda$ (see Eq. 2), need for unconditional layers in the discriminator, and distinction from ”coarse-to-fine” training strategies are presented in the Supp. C.
@@
Can any low-resolution be the starting point for high-res generations? In general, our observations indicate that high-resolution images are guided, yet not entirely restricted by the low-resolution images, as shown in Fig. 4. Each row in this figure presents a set of different high-resolution images that are generated from the same unconditional low-resolution image. This illustration reveals that high-resolution images can substantially differ from low-resolution ones, both in terms of color and texture. This suggests that subsequent conditional blocks are capable of significant modification in the background or foreground.
Importance of FS metrics: comparison against unconditional training. We also include results from training an unconditional StyleGAN2-ADA model [17] in Table 1, referred to as StyleGAN2-ADA UnCond. The unconditional model generates samples that follow the training distribution, which is mainly dominated by head classes. This bias favors unconditional learning in terms of the FID/KID metrics, which do not consider the skewness in the data distribution. Nevertheless, it can be observed that UnCond. baselines performed significantly worse in terms of few-shot metrics FID-FS/KID-FS. This suggests that relying solely on FID/KID can be misleading. We strongly recommend including few-shot metrics, i.e., FID-FS and KID-FS, when evaluating cGANs on long-tailed datasets.
CIFAR10-LT In this benchmark, our method is compared with various baselines with imbalance ratios $\rho$ of $50$ and $100$. This indicates that the least frequent tail class includes only $100$ and $50$ training instances, respectively. Table. 3 quantitatively assesses the quality of the generated samples by our method and the baselines. All the baselines use the StyleGAN2-ADA [17] architecture. Our method consistently outperforms the baselines across all metrics. This demonstrates that beyond improving generative learning across all classes (FID & KID), UTLO can also notably improve the learning from tail classes (FID-FS & KID-FS).
@@
Knowledge-sharing at Low Resolutions. As explained in Sec. 3, our proposed generator builds on top of low-resolution (e.g., $8\times 8$) unconditionally trained images which are subsequently used to generate conditional images at higher resolutions (e.g., $32\times 32$). Several low-resolution images are shown in Fig. 6 along with their high-resolution conditional images generated for the head and tail classes of the dataset. It can be seen that conditional images generated from the tail classes evidently share certain features from the head classes. This demonstrates that knowledge sharing at low resolutions is an effective approach for infusing information from head classes to tail ones. A quantitative analysis of knowledge-sharing is presented in Supp. D. For additional qualitative examples see Supp. H.
Figure 6: Knowledge sharing from head to tail classes in CIFAR10-LT dataset ($\rho=100$) using UTLO. The conditional images generated from the head (middle columns) and tail (right columns) classes share and are built on top of the same low-resolution (unconditional) images (left columns). Low-resolution images ($8\times 8$) are upsampled to that of CIFAR10-LT ($32\times 32$) for improved visualization.
CIFAR100-LT In addition to CIFAR10-LT, we evaluated our method and various baselines on the more challenging CIFAR100-LT dataset, where the tail classes include as few as five training instances. The quantitative results, shown in Table. 4, reveal that our method outperforms the baselines in a benchmark where a large number of classes with high diversity are present. In Fig. 2, we demonstrate the FID-FS curve during the course of the training along with the generated samples from the tail classes of the CIFAR100-LT dataset. This shows the effectiveness of our proposed method in addressing mode collapse while regularization methods diverge and require early stopping. Further visual comparisons of our proposed method against the baselines for CIFAR100-LT can be found in Supp. H.
@@
LSUN5-LT We include LSUN5-LT as a higher resolution ($128\times 128$) and highly-imbalanced ($\rho=1000$) benchmark. We use Projected GAN with FastGAN[26] generator and DA [40]) as the baseline model. This is a different generator and discriminator design compared to StyleGAN2 along with a different augmentation method. In the conditional implementation of FastGAN, class-specific information is injected using class-conditional batch normalization, different from the style mapping network utilized in StyleGAN2. UTLO can be readily adapted to this design by using standard batch normalization layers at low resolutions and utilizing class-conditional batch normalization only at higher resolutions. However, it should be noted that methods such as [37] are restricted to StyleGANs and cannot be applied in this context.
Table. 5 compares the results obtained from our method against the baseline and with the addition of GSR [36]. Our method consistently surpasses the baselines by around $30\%$ across all metrics. As previously shown for CIFAR10-LT (Table 3), UTLO not only improved on the tail classes but also significantly improved the dataset-wide FID/KID due to the high imbalance ratio ($\rho=1000$).
@@ 5 Conclusion
In this paper, we proposed UTLO, a novel framework designed to improve the training of cGANs on long-tailed data. Inspired by the observation that head and tail classes often have more similarities at lower resolutions, our method facilitates knowledge sharing from head to tail classes using unconditional training at lower resolutions. The proposed method enriches the limited training distribution of the tail classes and effectively addresses mode collapse, leading to significant improvement in image generation for tail classes. We have also introduced the FID-FS/KID-FS metric, an adaptation of widely-used GAN metrics, specifically tailored for tail classes. We hope that our findings are useful for future work in long-tail learning.
Acknowledgements
This work is partially supported by the National Science Foundation (NSF) under grant 1751412 and 1927564.
@@ Appendix A Datasets
  • CIFAR10/100: The original CIFAR10 and CIFAR100 datasets consist of 50,000 training images at $32\times 32$ resolution with an equal number of images present in 10 and 100 classes, respectively. An imbalanced version of these datasets referred to as CIFAR10/100-LT, is widely used in the long-tail recognition literature [49]. We follow [5] to generate an exponentially long-tailed version with an imbalance ratio $\rho=\{50,100\}$ for CIFAR10-LT and $\rho=100$ for CIFAR100-LT. When calculating few-shot metrics, we follow [4] and take classes 6-9 in CIFAR10-LT and 70-99 in CIFAR100-LT as the few-shot subset in our evaluation.
@@
Across all datasets, we sort image categories in the decreasing order of their size, i.e., class index 0 containing the most training images and so on.
@@ Appendix B Effects of Weighted Sampling
where $n_{c}$ is the number of samples in class $c$, $\rho=\max_{c}\{n_{c}\}/\min_{c}\{n_{c}\}$ is the imbalance ratio, and $\beta\in[0,1]$ is a hyperparameter that damps the sampling imbalance. We illustrate this in Fig. 7 for CIFAR100-LT with $\rho=100$. When $\beta=0$, the sampling follows the original data distribution. As $\beta$ increases, the effective imbalance ratio $\varrho$ among the classes decreases. At $\beta=1$, head and tail classes have the same probability of being drawn during training, i.e. $\varrho=1$. While WS might help with more calibrated training, it does not promote knowledge sharing among classes.
Figure 7: WS CIFAR100-LT $(\rho=100)$. When $\beta=0$, the sampling follows the original data distribution. As $\beta$ increases, the effective imbalance ratio $\varrho$ among the classes decreases. At $\beta=1$, head and tail classes have the same probability of being drawn during training, i.e. $\varrho=1$.
Table. 6 presents a comparison of the results obtained from the StyleGAN2-ADA baseline and the use of weighted sampling (WS). Our experiments indicate that WS improves the FID and KID metrics compared to the baseline. However, it does not enhance tail performance in terms of FID-FS and KID-FS, indicating that balancing alone is insufficient when there are very limited training examples in the tail. We also noticed that adding weighted sampling will boost overfitting. Fig. 8 illustrates the training FID-FS curves for the WS methods with different $\beta$ values, compared against our method and the baseline. For both CIFAR10-LT ($\rho=100$) and CIFAR100-LT ($\rho=100$) datasets, WS methods exhibit overfitting behavior. As $\beta$ increases, this becomes more pronounced, and training becomes unstable when $\beta=1$ across both datasets. While WS did not show any improvements for CIFAR100-LT, we found it to demonstrate relatively better performance than the baseline at the early stages of training before overfitting sets in. While our experiments on the role of WS in training cGANs in the long-tail setup are informative, we believe that reaching a comprehensive conclusion requires further analysis.
@@ Appendix C Ablation Study
Studying the results obtained in Table. 7, we observe that resolutions of $8\times 8$ and $16$ achieve relatively close performance. On the other hand, the performance degrades when layers up to $32\times 32$ are trained unconditionally. This is anticipated as the AnimalFaces-LT is at $64\times 64$, leaving only one layer to learn the class-conditional information which is shown to be insufficient.
To better understand the role of the intermediate low resolution selected during training, we show the unconditional low-res images $\hat{{\bm{x}}}_{l}$ from the intermediate layers (see Fig.3 in the main paper) for different ablated resolutions in Fig. 9. The low-res images are upsampled to the same size for better visual comparison. As $res{\text{uc}}$ increases, more details are introduced to the unconditionally trained intermediate images. For the AnimalFaces-LT dataset, resolution $32\times 32$ already includes fine and definite details, making it challenging to generate (tail) class-specific changes in a single layer before reaching the output resolution of $64\times 64$. We demonstrate this in the bottom of Fig. 9 the final class-conditional images at $64\times 64$ resolution from the tail class bear are shown. When $res_{\text{uc}}$ is higher, the final output changes minimally from the lower-resolution images $\hat{{\bm{x}}}l$. On the other hand, at lower $res{\text{uc}}$, the level of change is higher.
@@
Need for unconditional layers in the discriminator ($\mathcal{L}_{uc}$) and end-to-end joint training with both unconditional and conditional objectives. In addition to the unconditional layers in the generator, we have found that unconditional layers should be explicitly present in the discriminator. Table 9 provides the ablation results where the unconditional layers are removed from UTLO (i.e., w/o $\mathcal{L}_{uc}$). This shows significantly worse performance, indicating the need for explicit unconditional discriminator on the low resolution.
Further, to demonstrate how the proposed method differs from other training strategies that promote coarse-to-fine learning, e.g. progressive training [41], we carefully design experiments to compare our proposed method against progressive training. We follow the progressive strategy in StyleGAN-XL [41], starting with training a stem at a low resolution. After training the stem, the training of higher-resolution layers is followed. For a more comprehensive analysis, we experimented with two stems: an unconditional stem and a conditional one.
Firstly, we observed the conditional stem exhibited mode collapse early in the training. For training the subsequent higher-resolution layers, we picked the best stem before the mode collapse. This training strategy yielded considerably worse results as shown in Table 9 (last row). Conversely, the unconditional low-resolution stem did not experience mode collapse. Indeed, Table 9 shows that using unconditional stem (second-to-last row) improved over the baseline. However, it was still significantly worse than UTLO, showing the benefit of our design: end-to-end joint training with both unconditional and conditional objectives.
Table 9:
×$\times 1000$
StyleGAN2-ADA [17] 51.451.4{51.4}51.4 $87.1$ 24.724.7{24.7}24.7 $35.9$
+ UTLO (Ours) $\bm{26.2}$ $\underline{48.4}$ $\underline{12.6}$ $\underline{19.6}$
+ UTLO w/o $\mathcal{L}_{uc}$ 63.7563.75{63.75}63.75 $87.35$ 35.0835.08{35.08}35.08 $35.78$
+ Progressive [41]: Uncond. Stem 37.7337.73{37.73}37.73 $64.57$ 21.7121.71{21.71}21.71 $31.53$
+ Progressive [41]: Cond. Stem 98.3598.35{98.35}98.35 $150.86$ 73.9773.97{73.97}73.97 $102.20$